# When Do Neural Networks Learn World Models?

### v3

@@ Abstract
Humans develop world models that capture the underlying generation process of data. Whether neural networks can learn similar world models remains an open problem. In this work, we present the first theoretical results for this problem, showing that in a multi-task setting, models with a low-degree bias provably recover latent data-generating variables under mild assumptions—even if proxy tasks involve complex, non-linear functions of the latents. However, such recovery is sensitive to model architecture. Our analysis leverages Boolean models of task solutions via the Fourier-Walsh transform and introduces new techniques for analyzing invertible Boolean transforms, which may be of independent interest. We illustrate the algorithmic implications of our results and connect them to related research areas, including self-supervised learning, out-of-distribution generalization, and the linear representation hypothesis in large language models.
@@ 1 Introduction
    In Section 3, we introduce complexity measures based on a notion of realization degree, offering an approach for analyzing the impact of low-complexity bias on world model learning.
@@
    In Section 4, we present the first theoretial results on learning world models in the context of training on proxy tasks using observed data. We identify two critical factors for world model learning: (i) a multi-task setting; (ii) the low-complexity bias, instantiated by a low-degree bias of the model and a low-degree task distribution. Together, these factors ensure the identifiability of latent data-generating variables. Moreover, we show the provable benefits of learning world models in an out-of-distribution generalization setting (Abbe et al., 2023) and study the impact of model architecture under a notion of basis compatibility (see Figure 1 for a graphical summary of our results). Technically, our analysis relies on analyzing the degree properties of Boolean functions composed with invertible transforms, which may be of independent interest.
@@ 2 Formulation of Learning World Models
Lemma 2.4 indicates that we can construct new random variables $\Phi({\mathbf{x}})={\mathbf{z}}^{\prime}$ that have the same distribution $p$ as true latent variables $\psi^{-1}({\mathbf{x}})={\mathbf{z}}$ (thus fitting the observed variables ${\mathbf{x}}$ equally well) by first transforming ${\mathbf{z}}$ to standard Gaussian variables, applying any orthogonal transform, and then inverting the transform (note that standard Gaussian distributions are invariant to orthogonal transforms). Applying this result to our context, we conclude that without additional assumptions, learning world models in the sense of Definition 2.2 is not possible with a simple ${\mathcal{T}}$ by looking at ${\mathbf{x}}$ alone.
@@ 3 Complexity Measures
Computers encode every object, including the observed data and variables learned by neural networks, in bit strings. Leveraging this fact, we can assume without loss of generality that both ${\mathbf{x}}$ and ${\mathbf{z}}$ in Definition 2.1 are Boolean (after some encoding). Formally, in the remainder of the paper we will let ${\mathcal{X}}\subseteq\{-1,1\}^{m}$ and ${\mathcal{Z}}=\{-1,1\}^{d}$. Since $\psi$ is invertible, we have $m\geq d$. In practice, we expect $m\gg d$ for complex and high-dimensional data. Unless otherwise mentioned, we will also assume $\mathrm{supp}(p)={\mathcal{Z}}$ to ensure that all elements in ${\mathcal{Z}}$ can be sampled with positive probabilities.
A direct consequence of the Boolean modeling of variables is that all functions involved are Boolean functions. Notably, this itself does not resolve the non-identifiability issue present in the continuous case, as functions with Boolean inputs and outputs can still exhibit arbitrary nonlinearity. Instead, the primary advantage of this approach lies in the useful machinery it offers for defining functional complexity through the Fourier-Walsh transform:
@@
The Fourier-Walsh transform shows that every Boolean function can be represented as a linear combination of parity functions $\chi_{S}$ that capture all non-linear relationship between inputs and outputs. In fact, it can be shown that parity functions are a basis of the vector space ${\mathcal{F}}^{n}\vcentcolon=\{f:\{\pm 1\}^{n}\to{\mathbb{R}}\}$ of $n$-dimensional Boolean functions (see Section D for more details). By contrast, the space of arbitrary continuous functions does not have a similar, theory-friendly basis. See Section B for more discussion on the use of Boolean functions and complexity measures.
@@
By Definition 3.5, flat realizations with the low-degree bias satisfy $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$. Analyzing the impact of representations is more involved; the main intuition behind Definition 3.6 is that a representation $\Phi$ only makes a task $h$ “simpler” if the conditional degree $\mathrm{deg}(h\mid\Phi)>0$, i.e., solving the task on top of $\Phi$ yields a smaller degree compared to low-degree flat realizations. This notion will be explored further in the next section.
@@ 4 Theoretical Analysis
This section presents our main theoretical results. We first study how the low-degree bias introduced in Section 3 drives representation learning (Section 4.1). We then present our main results on learning world models, showing sufficient conditions for identifying latent data-generating variables (Section 4.2). Next, we show provable benefits of learning world models in an out-of-distribution generalization setting (Section 4.3). We conclude this section by a preliminary analysis on how the model architecture impacts world model learning (Section 4.4). All proofs are deferred to Section E. Finally, Section C presents numerical experiments that substantiate our theoretical results.
@@ 5.1 Polynomial Extrapolation
One may wonder that in this task, we still need to know the basis of the task a priori to achieve basis compatibility. However, we emphasize that even without prior knowledge, we can still use different activation functions simultaneously and rely on the low-complexity bias of neural networks to self-adaptively select functions that are the most compatible with the task, as empirically shown in our experiments. We thus conjecture that this approach could be quite universal. Indeed, in the next section we show that exactly the same approach can also bring benefits in a distinct scenario.
@@ Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under STI 2030-Major Projects 2021ZD0200300, and in part by the National Key Research and Development Program of China No. 2024YDLN0006, and in part by the National Natural Science Foundation of China under Grant 62176133.
@@ Appendix A Related Work
A growing body of research introduces complexity measures tailored for neural networks trained via (stochastic) gradient descent (Bartlett et al., 2017; Jacot et al., 2018; Zhou et al., 2019; Lotfi et al., 2022; Chatterjee & Sudijono, 2024), often leveraging them for generalization analysis. However, these measures inherently depend on the neural network parameterization, rather than capturing function complexity independently of specific parameterizations. Some studies propose alternative complexity metrics inspired by Kolmogorov complexity in the machine learning context (Xu et al., 2020; Liu et al., 2023), but these metrics are typically problem-specific and do not enable a direct complexity analysis in the function space, unlike the approach we take in this work. See Section B for more discussion on some commonly-used complexity measures.
Appendix B Discussion on Complexity Measures
Appendix C Numerical Experiments
  • Identifying data-generating variables benefits OOD generalization: Figure 5 shows that in a Hamming sampling setting as in Theorem 4.9, MLPs with a representation $\Phi$ that identifies true data-generating variables achieve superior OOD generalization compared to MLPs without such a representation.
Please refer to the figures for more details and discussion.
$x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $x_{8}$ $x_{9}$ $x_{10}$
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
(a) $2$-degree tasks
(b) $3$-degree tasks
(c) $4$-degree tasks
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
Figure 5: Identifying data-generating variables benefits out-of-distribution generalization: For the same tasks considered in Figure 3 with $r=2$, SGD-trained MLPs on top of an oracle representation that identifies true data-generating variables (denoted by “w/ WM” in the legends) achieve out-of-distribution (OOD) generalization. By contrast, SGD-trained MLPs without the representation (denoted by “w/o WM” in the legends) fail. This corroborates Theorem 4.9 in the main text. We separately train an MLP for each task and for the setting with/without the oracle representation.
Appendix D Preliminaries on Boolean Function Analysis
@@ E.1 Technical Lemmas
for every ${\bm{x}}\in{\mathcal{X}}^{\prime}\vcentcolon=\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$—if this does not hold, then we have $|{\mathcal{X}}^{\prime}|\geq 2^{k^{\prime}}>2^{k}\geq\sum_{i=0}^{r}{d\choose r}$, which contradicts $|{\mathcal{X}}^{\prime}|=\sum_{i=0}^{r}{d\choose r}$. By equation (64), we can replace every degree-$k^{\prime}$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$k^{\prime}-1$ polynomial $-\sum_{S^{\prime}\subset S,|S^{\prime}|\leq k^{\prime}-1}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. Iteratively using this replacement in the Fourier-Walsh transform of $h$ gives the desired result. ∎
@@ G.1 Polynomial Extrapolation
    Xu et al. (2021) replaces all ReLU functions with quadratic activation functions, while we only replace half of ReLU functions with quadratic activation functions and identity activation functions. This difference is important in scenarios where we do not know the exact structural form of target functions—note that our method can be viewed as an “ensemble” of different activation functions, which enables the neural network to adaptively select activation functions that are the most compatible with the task as shown by our empirical results.

### v4

@@ Abstract
Humans develop world models that capture the underlying generation process of data. Whether neural networks can learn similar world models remains an open problem. In this work, we present the first theoretical results for this problem, showing that in a multi-task setting, models with a low-degree bias provably recover latent data-generating variables under mild assumptions–even if proxy tasks involve complex, non-linear functions of the latents. However, such recovery is sensitive to model architecture. Our analysis leverages Boolean models of task solutions via the Fourier-Walsh transform and introduces new techniques for analyzing invertible Boolean transforms, which may be of independent interest. We illustrate the algorithmic implications of our results and connect them to related research areas, including self-supervised learning, out-of-distribution generalization, and the linear representation hypothesis in large language models.
@@ 1 Introduction
    In Section 3, we introduce our Boolean function formulation and the corresponding complexity measures based on a notion of realization degree, offering a tractable approach for analyzing the impact of low-complexity bias on world model learning.
@@
    In Section 4, we present the first theoretial results on learning world models in the context of training on proxy tasks using observed data. We identify two critical factors for world model learning: (i) a multi-task learning setting; (ii) the low-complexity bias, instantiated by a low-degree bias of the model and a low-degree task distribution. Together, these factors ensure the identifiability of latent data-generating variables. Moreover, we show the provable benefits of learning world models in an out-of-distribution generalization setting (Abbe et al., 2023) and study the impact of model architecture under a notion of basis compatibility (see Figure 1 for a graphical summary of our results). Technically, our analysis relies on analyzing the degree properties of Boolean functions composed with invertible transforms, which may be of independent interest.
@@ 2 Formulation of Learning World Models
Lemma 2.4 indicates that we can construct new random variables $\Phi({\mathbf{x}})={\mathbf{z}}^{\prime}$ that have the same distribution $p$ as true latent variables $\psi^{-1}({\mathbf{x}})={\mathbf{z}}$ (thus fitting the observed variables ${\mathbf{x}}$ equally well) by first transforming ${\mathbf{z}}$ to standard Gaussian variables, applying any orthogonal transform, and then inverting the transform (note that standard Gaussian distributions are invariant to orthogonal transforms). Applying this result to our context, we conclude that without additional assumptions, learning world models in the sense of Definition 2.2 is not possible with a simple ${\mathcal{T}}$ by looking at the observed data ${\mathbf{x}}$ alone.
@@ 3 Complexity Measures
Computers encode every object, including the observed data and variables learned by neural networks, in bit strings. Motivated by this fact, we could assume without loss of generality that both ${\mathbf{x}}$ and ${\mathbf{z}}$ in Definition 2.1 are Boolean (after some proper encoding). Formally, in the rest of the paper we will let ${\mathcal{X}}\subseteq\{-1,1\}^{m}$ and ${\mathcal{Z}}=\{-1,1\}^{d}$. Since $\psi$ is invertible, we have $m\geq d$. Unless otherwise mentioned, we will also assume $\mathrm{supp}(p)={\mathcal{Z}}$ to ensure that all elements in ${\mathcal{Z}}$ can be sampled with non-zero probabilities.
A direct consequence of the Boolean modeling of variables is that all functions involved are Boolean functions (see Section B for a brief introduction). Notably, this itself does not resolve the non-identifiability issue stated in Section 2, as functions with Boolean inputs and outputs can still exhibit arbitrary nonlinearity. Instead, the primary advantage of this treatment is the useful machinery it provides for defining functional complexity, offered by the Fourier-Walsh transform of Boolean functions:
@@
The Fourier-Walsh transform shows that every Boolean function can be represented as a linear combination of parity functions $\chi_{S}$ that capture all non-linear relationship between inputs and outputs. It can also be shown that parity functions are a basis of the vector space ${\mathcal{F}}^{n}\vcentcolon=\{f:\{\pm 1\}^{n}\to{\mathbb{R}}\}$ of $n$-dimensional Boolean functions (see Section B for more details). By contrast, the space of arbitrary continuous functions does not have a similar, theory-friendly basis. See Section C for more discussion on the use of Boolean functions and complexity measures.
@@
By Definition 3.5, flat realizations with low-degree bias satisfy $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$. Analyzing the impact of representations is more involved; the main intuition behind Definition 3.6 is that a representation $\Phi$ only makes a task $h$ “simpler” if the conditional degree $\mathrm{deg}(h\mid\Phi)>0$, i.e., solving the task on top of $\Phi$ yields a smaller degree compared to low-degree flat realizations. This notion will be explored further in the next section.
@@ 4 Theoretical Analysis
This section presents our main theoretical results. We first study how the low-degree bias introduced in Section 3 drives representation learning (Section 4.1). We then present our main results on learning world models, showing sufficient conditions for identifying latent data-generating variables (Section 4.2). Next, we show provable benefits of learning world models in an out-of-distribution generalization setting (Section 4.3). We conclude this section by a preliminary analysis on how the model architecture impacts world model learning (Section 4.4). All proofs are deferred to Section E. Finally, Section D presents numerical experiments that substantiate our theoretical results.
@@ 5.1 Polynomial Extrapolation
One may wonder that in this task, we still need to know the basis of the task a priori to achieve basis compatibility. However, we emphasize that even without prior knowledge, we can still use different activation functions simultaneously and rely on the low-complexity bias of neural networks to self-adaptively select functions that are the most compatible with the task, as empirically shown in our experiments. We thus conjecture that this approach could be quite universal. Indeed, in the next section we will show that exactly the same approach can also bring benefits in a distinct scenario.
5.2 Learning Physical Laws
@@ Acknowledgements
This work was supported in part by the National Key Research and Development Program of China under STI 2030–Major Projects 2021ZD0200300, and in part by the National Key Research and Development Program of China No. 2024YDLN0006, and in part by the National Natural Science Foundation of China under Grant 62176133.
@@ Appendix A Related Work
A growing body of research introduces complexity measures tailored for neural networks trained via (stochastic) gradient descent (Bartlett et al., 2017; Jacot et al., 2018; Zhou et al., 2019; Lotfi et al., 2022; Chatterjee & Sudijono, 2024), often leveraging them for generalization analysis. However, these measures inherently depend on the neural network parameterization, rather than capturing function complexity independently of specific parameterizations. Some studies propose alternative complexity metrics inspired by Kolmogorov complexity in the machine learning context (Xu et al., 2020; Liu et al., 2023), but these metrics are typically problem-specific and do not enable a direct complexity analysis in the function space, unlike the approach we take in this work. See Section C for more discussion on some commonly-used complexity measures.
Appendix B Preliminaries on Boolean Function Analysis
@@
Appendix C Discussion on Complexity Measures
Appendix D Numerical Experiments
  • Identifying data-generating variables benefits out-of-distribution (OOD) generalization: Figure 5 shows that in a Hamming sampling setting as in Theorem 4.9, MLPs with a representation $\Phi$ that identifies true data-generating variables achieve superior OOD generalization compared to MLPs without such a representation.
Please refer to the figures for more details and discussion.
$x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $x_{6}$ $x_{7}$ $x_{8}$ $x_{9}$ $x_{10}$
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
(a) $2$-degree tasks
(b) $3$-degree tasks
(c) $4$-degree tasks
(a) Task 1: $f({\bm{x}})=x_{1}x_{4}x_{5}$
(b) Task 2: $f({\bm{x}})=x_{1}x_{3}+0.5x_{1}x_{5}x_{6}$
Figure 5: Identifying data-generating variables benefits out-of-distribution (OOD) generalization: For the same tasks considered in Figure 3 with $r=2$, SGD-trained MLPs on top of an oracle representation that identifies true data-generating variables (denoted by “w/ WM” in the legends) achieve OOD generalization. By contrast, SGD-trained MLPs without the representation (denoted by “w/o WM” in the legends) fail. This corroborates Theorem 4.9 in the main text. We separately train an MLP for each task and for the setting with/without the oracle representation.
@@ E.1 Technical Lemmas
for every ${\bm{x}}\in{\mathcal{X}}^{\prime}\vcentcolon=\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$–if this does not hold, then we have $|{\mathcal{X}}^{\prime}|\geq 2^{k^{\prime}}>2^{k}\geq\sum_{i=0}^{r}{d\choose r}$, which contradicts $|{\mathcal{X}}^{\prime}|=\sum_{i=0}^{r}{d\choose r}$. By equation (64), we can replace every degree-$k^{\prime}$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$k^{\prime}-1$ polynomial $-\sum_{S^{\prime}\subset S,|S^{\prime}|\leq k^{\prime}-1}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. Iteratively using this replacement in the Fourier-Walsh transform of $h$ gives the desired result. ∎
@@ G.1 Polynomial Extrapolation
    Xu et al. (2021) replaces all ReLU functions with quadratic activation functions, while we only replace half of ReLU functions with quadratic activation functions and identity activation functions. This difference is important in scenarios where we do not know the exact structural form of target functions–note that our method can be viewed as an “ensemble” of different activation functions, which enables the neural network to adaptively select activation functions that are the most compatible with the task as shown by our empirical results.